### v1

# Learning from Mistakes: Enhancing Harmful Meme Detection via Misjudgment Risk Patterns

Abstract.
Internet memes have emerged as a popular multimodal medium, yet they are increasingly weaponized to convey harmful opinions through subtle rhetorical devices like irony and metaphor. Existing detection approaches, including MLLM-based techniques, struggle with these implicit expressions, leading to frequent misjudgments. This paper introduces PatMD, a novel approach that improves harmful meme detection by learning from and proactively mitigating these potential misjudgment risks. Our core idea is to move beyond superficial content-level matching and instead identify the underlying misjudgment risk patterns, proactively guiding the MLLMs to avoid known misjudgment pitfalls. We first construct a knowledge base where each meme is deconstructed into a misjudgment risk pattern explaining why it might be misjudged, either overlooking harmful undertones (false negative) or overinterpreting benign content (false positive). For a given target meme, PatMD retrieves relevant patterns and utilizes them to dynamically guide the MLLM’s reasoning. Experiments on a benchmark of 6,626 memes across 5 harmful detection tasks show that PatMD outperforms state-of-the-art baselines, achieving an average of 8.30% improvement in F1-score and 7.71% improvement in accuracy, demonstrating strong generalizability and improved detection capability of harmful memes.
Disclaimer: This paper contains content for demonstration purposes that may be disturbing to some readers.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; $X$, 2025; XXisbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
With the widespread adoption of the Internet, memes have emerged as one of the most vibrant cultural phenomena of the digital age. These multimodal information units, which ingeniously combine images and text, leverage their humorous, satirical, and highly viral nature to proliferate across social media platforms. This phenomenon has profoundly reshaped the modes of information production and consumption. Beyond being a medium for entertainment and social interaction, memes have evolved into a potent tool for expressing viewpoints, fostering community cohesion, and even influencing public opinion (Lin et al., 2024). However, this powerful influence is a double-edged sword. Memes have also become a vessel for packaging and amplifying hate speech, reinforcing biases, and disseminating implicit discrimination and misinformation (Sharma et al., 2022), posing a severe threat to the health of the online platforms.
To safeguard online platforms from the pervasive negative impact of harmful memes, developing automated detection approaches has emerged as a critical area of research (Kiela et al., 2020). Initial efforts in this domain largely adopted a feature-fusion paradigm, first processing the image and text through independent encoders and then combining their representations to identify harmful content (Suryawanshi et al., 2020; Pramanick et al., 2021). Subsequently, the field shifted towards fine-tuning large pre-trained multimodal models, which jointly model visual and textual inputs to achieve superior cross-modal alignment (Lu et al., 2019; Chen et al., 2020). More recently, to address the need for deeper semantic reasoning, research has pivoted towards leveraging the capabilities of Multimodal Large Language Models (MLLMs), either by generating explicit reasoning chains (Hee et al., 2022; Lin et al., 2023) or by retrieving similar samples to augment the decision-making process (Liu et al., 2025).
Figure 1. The motivation example of harmful meme detection with misjudgment risk patterns
While MLLMs demonstrate a strong capacity for content comprehension, they exhibit notable limitations in detecting harmful memes. They may correctly identify all elements within a meme, yet lack the conceptual reasoning capability necessary to navigate complex contextual cues—particularly when confronted with sarcasm, metaphor, or implicit stereotypes. This deficiency often leads to erroneous misjudgments. As shown in Figure 1, the MLLMs accurately identify the elements in the memes, such as humans and monkeys, but they cannot capture the underlying decision logic behind harmful intent, i.e., why & how the combination of elements causes racially hateful memes.
As previously mentioned, due to the prevalent use of rhetorical devices such as irony and analogy in harmful memes, MLLMs frequently commit misjudgments in specific, systematic ways. Our core insight is to preemptively address this issue by abstracting these failure modes into misjudgment risk patterns. These patterns serve as explicit guidance, cautioning the MLLM to guard against these common pitfalls during its final decision-making. As such, our work presents a paradigm shift: moving from a purely content-centric detection to a model decision-centric one. By focusing not only on whether a meme is harmful but also on how the model reasons about it and why that reasoning may fail, we directly mitigate characteristic cognitive and interpretive limitations of MLLMs in meme analysis. As illustrated in Figure 1, we derive the misjudgment risk pattern from similar memes, indicating that the racially hateful information comes from the incorrect human and non-human comparison and causes a deliberate dehumanization. With the help of this pattern, MLLMs go beyond the separated human and animal elements and capture the underlying decision logic from human-animal relations. The misjudgment is mitigated, and MLLMs successfully identify the racially hateful memes.
In this paper, we propose PatMD, an automated approach that leverages misjudgment risk patterns to improve harmful meme detection. PatMD constructs a knowledge base by diagnostically decomposing memes to archive past MLLM misjudgments. For any new meme, it retrieves these misjudgment risk patterns and injects them the MLLM’s reasoning process, proactively steer the analysis away from known pitfalls. PatMD consists of three stages: In misjudgment risk pattern elicitation stage, historical memes are mapped into a hierarchical harm tree, capturing multi-granular harmful concepts, from which corresponding misjudgment risk patterns are derived. Then risk-aware pattern retrieval stage identifies semantically similar memes based on structural similarity within the hierarchical harm tree, and fetches their associated risk patterns. In pattern-augmented reasoning for the detection stage, the retrieved patterns are integrated into a structured prompt, guiding the MLLMs to avoid known misjudgment pitfalls.
To evaluate the performance of PatMD, we conduct extensive experiments on a benchmark with 6,626 memes spanning 5 harmful detection tasks. The results demonstrate that PatMD have substantial and consistent improvements against four specialized baselines and six general-purpose MLLMs, with an average improvement of 8.30% in F1-score and 7.71% in accuracy. The ablation study reveals the necessity of each component in our approach design. Further analysis reveals that the strength of PatMD lies in its ability to steer MLLMs toward proactive misjudgment prevention by focusing on the underlying logic of harm manifestation, rather than relying on superficial content analysis. This misjudgment-aware reasoning paradigm enables the model to perform more accurate and nuanced judgment, effectively reducing prevalent misclassification errors.
The main contributions of this paper are summarized as follows:
  • We propose PatMD, a novel approach that improves harmful meme detection by learning from potential misjudgment risk patterns of historical memes. This presents a paradigm shift for harmful meme detection: moving from a purely content-centric detection to a model decision-centric one, i.e., proactively steering the MLLM’s reasoning away from known misjudgment pitfalls.
  • We conduct comprehensive experiments on a benchmark covering five harmful detection tasks, comparing with four specialized baselines and six general-purpose MLLMs. The results demonstrate its effectiveness, achieving an average of 8.30% improvement in F1-score and 7.71% improvement in accuracy.
  • Our code and dataset are available  [^1] https://anonymous.4open.science/r/PatMD-61BF for replication.
2. Related Works
Harmful meme detection has developed into an important direction in multimodal research and has continued to advance with the introduction of large-scale benchmark datasets (Cai et al., 2019; Kiela et al., 2020; Fersini et al., 2022; Lin et al., 2024). Early studies mainly focused on feature fusion, typically adopting a two-stream architecture that extracts visual and textual representations through independent encoders and then leverages attention mechanisms for cross-modal interaction and alignment (Suryawanshi et al., 2020; Pramanick et al., 2021; Lee et al., 2021). Subsequently, research shifted towards shared Transformer-based intermediate fusion and fine-tuning of pretrained multimodal models, both of which jointly model image regions and text sequences to enhance cross-modal alignment and discrimination (Kiela et al., 2019; Lu et al., 2019; Li et al., 2019; Chen et al., 2020). With the introduction of the Hateful Memes Challenge (Kiela et al., 2020), a variety of ensemble-based and contrastive reweighting approaches emerged (Muennighoff, 2020; Sandulescu, 2020; Lippe et al., 2020), further improving their generalizability. Other researchers explore prompt-based methods. Some work in this area (Cao et al., 2022; Ji et al., 2023; Cao et al., 2023; Ji et al., 2024) focused on concatenating the meme’s text with extracted image captions to fine-tune pre-trained language models(Liu et al., 2019; Chung et al., 2024) for the detection task. Additionally, recent work (Lin et al., 2025) has employed an interactive, multi-agent framework for deeper visual understanding. This framework simulates a discussion where agents query a vision expert to generate a synthesized analysis.
Figure 2. The overview of PatMD:
Large Language Models (LLMs) have demonstrated exceptional capabilities in complex reasoning (Brown et al., 2020; Chowdhery et al., 2023), such as generating intermediate reasoning steps via Chain-of-Thought (CoT) prompting to enhance judgment accuracy (Kojima et al., 2022; Wei et al., 2022; Zhang et al., 2022). Inspired by this, research in interpretable harmful meme detection has begun to leverage LLMs to generate explicit chains of reasoning to explain their rationale. These methods, such as constructing LLM debate frameworks (Hee et al., 2022) or distilling LLM’s reasoning knowledge into smaller models (Lin et al., 2023). To address challenges in low-resource generalization, researchers have explored few-shot enhanced methods (Cao et al., 2024; Huang et al., 2024). Inspired by this direction, Liu et al. introduced MIND (Liu et al., 2025), a framework for zero-shot harmful meme detection that integrates similar example retrieval and multi-agent collaboration. This aligns with the broader paradigm of Retrieval-Augmented Generation (RAG) (Lewis et al., 2020; Guu et al., 2020), which enriches an LLM’s input with external knowledge and has proven highly effective in knowledge-intensive tasks (Ram et al., 2023; Jiang et al., 2023; Asai et al., 2024; Wei et al., 2024). While the aforementioned approaches improve performance by enhancing content understanding or providing content examples, they remain fundamentally content-centric. When adversaries design novel memes, these methods may struggle to find relevant content for reference. Inspired by these lines of research, we propose a paradigm shift. Instead of focusing on ”what the content is,” our method focuses on ”why the model might misjudge.” We introduce an approach based on misjudgment risk patterns and integrate it into the LLM’s decision-making process. This model-decision-centric perspective allows it to identify and guard against specific reasoning paths that lead to errors.
3. Overview of PatMD
Our approach, PatMD, is designed to enhance the harmful meme detection capabilities of MLLMs. We shift the paradigm from reliance on the MLLM’s implicit knowledge to proactively guide it with explicit misjudgment risk patterns derived from historical memes. These patterns capture the specific manner in which the MLLM may misinterpret the multimodal content. During inference, relevant patterns are used to steer the MLLM’s reasoning process, proactively alerting it to potential pitfalls and thereby enabling more robust judgments. The implementation of PatMD is structured into three main stages, as demonstrated in Figure 2.
Misjudgment Risk Pattern Elicitation: We deconstruct historical memes into a hierarchical harm tree, which structurally represents harmful concepts through a macro-to-micro schema. This multi-level representation then serves as the foundation for generating the meme’s corresponding misjudgment risk pattern, which captures how its semantic and contextual composition may trigger specific reasoning failures in MLLMs.
Risk-aware Pattern Retrieval: For a target meme, this stage performs a two-step, risk-aware retrieval process. After gathering candidate memes based on coarse content similarity, it then reranks them using a hierarchical harm tree-based similarity measure to identify those with the most analogous harmful composition.
Pattern-Augmented Reasoning for Detection: Finally, the patterns from the retrieved memes are dynamically integrated into a structured prompt. This guides the MLLM to perform a well-reasoned analysis, safeguarding it against the misjudgments identified in the historical patterns.
3.1. Misjudgment Risk Pattern Elicitation
This Misjudgment Risk Pattern Elicitation stage forms the cornerstone of our approach, which provides the patterns that are used to steer the MLLM’s reasoning process. This stage first conducts the hierarchical harm deconstruction to systematically capture the diversified views of the harmful concepts in a meme, followed by the causal misjudgment attribution, aiming at deriving the misjudgment risk pattern.
3.1.1. Hierarchical Harm Deconstruction
To systematically capture the diversified view of harmful composition, we first perform the hierarchical harm deconstruction. This process employs a “Macro to Micro” schema to structurally dissect a meme’s harmful composition, delineating its underlying harmful logic layer by layer and moving beyond superficial features to reveal its core harmful structure. The results across all memes collectively form a Hierarchical Harm Tree $\mathcal{T}_{\text{H}}$, where each meme corresponds to a distinct structural harm instance within the tree.
Hierarchical Harm Tree Structure. The tree incorporates a root node $RNode$ and other tree nodes $TNode$. If the tree node is at the bottom level, it is considered a leaf node $LNodes$. The tree can be formalized as follows:
(1) $\mathcal{T}_{\text{H}}=\left\{\begin{array}[]{lr}RNode\rightarrow[TNode^{(1)}_{L_{1}},...,TNode^{(N)}_{L_{1}}]&\\ TNode_{L_{i}}^{(j)}\rightarrow[TNode^{(1)}_{L_{i+1}},...,TNode^{(N_{j})}_{L_{i+1}}],\\ \quad\text{where}\;i\in\{1,...,4\}\;\text{and}\;TNode_{L_{4}}^{(i)}=LNode\\ LNode_{L_{4}}^{(i)}\rightarrow\emptyset\end{array}\right$.
where the symbol $\rightarrow$ is the direction of the tree’s layer. For each tree node $TNode^{(j)}_{L_{i}}$, the subscript $L_{i}$ means the level of this node, and the superscript $(j)$ is the $j_{th}$ node of the $L_{i}$ level’s node, and $N_{j}$ is the number of its child nodes. Based on the harm expressions in memes, and inspired by the framework for harmful category classification in (Chen et al., 2025), we define a four-level tree $L_{1},...,L_{4}]$. The content of each layer’s node is defined as follows:
  • $L_{1}$ (Macro Category): Identifies the category of the content (i.e., Nationality, Gender, Religion, Race, Animal, Disability, Child Exploitation, Political, None)
  • $L_{2}$ (Core Subjects): Recognizes the specific groups, symbols, or figures explicitly targeted in the meme.
  • $L_{3}$ (Expressive Techniques): Analyzes the key methods used by the creator to convey the message (e.g., use of stereotypes, irony, direct assertion).
  • $L_{4}$ (Instantiation of Techniques): Describes the concrete manifestation of the aforementioned techniques in the meme.
Automated Tree Construction. Guided by the defined hierarchical structure, we employ an MLLM to automatically decompose every meme’s content into conceptual nodes at each level of the hierarchical harm tree.
(2) $\mathcal{T}_{\text{H}}^{(i)}=\text{MLLM}_{\text{gen}}(m_{i},\text{Template}(\mathcal{T}_{\text{H}}))$
where for a given meme $m_{i}$, the generation function $\text{MLLM}_{\text{gen}}$ fills the nodes of the tree template $\text{Template}(\mathcal{T}_{\text{H}})$ to produce the final instantiated tree $\mathcal{T}_{\text{H}}^{(i)}$ (See Appendix C Figure 7 for the detailed prompt templates). This process yields a structural harm instance in the hierarchical harm tree. For instance, as shown in Figure 3, the MLLM would identify the ‘Macro Category’ ($L_{1}$) as “Race”, the ‘Core Subject’ ($L_{2}$) as “African descent”, and so on.
Figure 3. Example of constructing the hierarchical harm tree and corresponding misjudgment risk pattern of the meme.
3.1.2. Causal Misjudgment Attribution
Building upon the deconstructed harm profile, we then derive the misjudgment risk pattern. We analyze how the identified harmful concepts could lead an MLLM to specific misjudgments (e.g., overlooking ironic cues or misinterpreting contextual references), and produce the misjudgment risk pattern for each meme. We posit that the processes for generating misjudgment risk patterns differ fundamentally between benign and hateful samples. Therefore, PatMD employs distinct analytical strategies for each:
For benign samples, PatMD infers the root cause of its false positive risk based on the structural harm information. It analyzes why the combination of its theme, entities, and expressive techniques might be misinterpreted or over-interpreted as harmful.
For harmful samples, PatMD reasons in reverse, determining how the meme’s compositional elements are subtly employed to encode or conceal its false negative risk—that is, its underlying harmful intent.
This dual-pathway analysis helps to address the possibility of both false positives and false negatives. The MLLM abstracts the harmful concepts into a concise, generalizable misjudgment risk pattern (See Appendix C Figure 9 for the detailed prompt templates).
After the above process, each sample $m_{i}=(\text{img}_{i},\text{txt}_{i})$ is endowed with a structural harm instance $\mathcal{T}_{i}$ within the hierarchical harm tree and a misjudgment risk pattern $\mathcal{P}_{i}$. To make these information retrievable, we also employ a pre-trained multimodal encoder $E$ (Günther et al., 2025) to generate embedding vectors for its image and text, $v^{\text{img}}_{i}=E(\text{img}_{i})$ and $v^{\text{txt}}_{i}=E(\text{txt}_{i})$, respectively. This stage ultimately produces a knowledge base where each meme is encoded as a tuple $(v^{\text{img}}_{i},v^{\text{txt}}_{i},\mathcal{T}_{i},\mathcal{P}_{i},\mathcal{G}_{i})$, recording its multimodal embeddings, structural harm instance, misjudgment risk pattern, and ground truth label (benign/harmful) $\mathcal{G}_{i}$.
3.2. Risk-aware Pattern Retrieval
To guide the follow-up detection of a target meme, PatMD performs a risk-aware pattern retrieval. We go beyond content similarity by seeking memes that potentially share a comparable risk profile, as reflected by the hierarchical harm tree. PatMD implements this by first gathering content-based candidates, then applying a hierarchical harm tree-based reranking to pinpoint the most structurally similar memes.
3.2.1. Bipartite Candidate Retrieval
The goal of the initial retrieval stage is to construct a diverse set of candidate samples. We employ a label-based bipartite retrieval strategy to include both benign and harmful samples.
Specifically, we use the target sample’s image vector $v_{\text{test}}^{\text{img}}$ and text vector $v_{\text{test}}^{\text{txt}}$ to separately query the subsets of samples corresponding to benign and harmful ground-truth labels. Within each labeled subset, we retrieve the top-$N$ most similar candidates from the image modality and the top-$N$ from the text modality. Subsequently, we take the union of all these retrieved candidates to form an initial candidate pool. This pool, containing samples that offer a balanced perspective on potential risks, is then passed to the reranking part for a more refined similarity assessment and re-ranking.
3.2.2. Hierarchical Harm Tree-based Reranking
The initial retrieval process, which relies on coarse vector similarity, may fail to adequately capture the nuanced manifestation of harmful concepts between the target and candidate memes. We utilize the structural matching of the hierarchical harm tree to choose more suitable memes.
We employ a top-down sequential matching mechanism that terminates at the first mismatched level. The matching sequence can start with a successful match at either the $L_{1}$ or $L_{2}$ level and proceed downwards. A match at the $L_{1}$ level requires an exact content match. For levels $L_{2}$ and below, a match is successful if the semantic similarity of their node descriptions exceeds a predefined threshold $\beta$.
For each successfully matched level ($L_{1}$, $L_{2}$, $L_{3}$, or $L_{4}$), a fixed bonus score $\gamma$ is added to the candidate’s similarity score, formulated as follows:
(3) $S_{\mathcal{T}}=\sum_{k=1}^{4}\gamma\prod_{i=1}^{k}M_{i}$
where $\gamma$ is the bonus for a match at level $L_{k}$. The match status $M_{i}\in\{0,1\}$ is determined by:
  • For $L_{1}$: $M_{1}=1$ if an exact content match occurs.
  • For $L_{i>1}$: $M_{i}=1$ if the semantic similarity of node content exceeds a threshold $\beta$.
The product term $\prod_{i=1}^{k}M_{i}$ ensures that the bonus $\gamma$ is awarded only if all levels from $L_{1}$ to $L_{k}$ have successfully matched, thus enforcing the top-down sequential mechanism. Consequently, the tree similarity score for a candidate, denoted as $S_{\mathcal{T}}$, is hierarchical bonuses accumulated from all continuously matched levels (top-down).
In addition, we also utilize a unified Multimodal Reranker [^2] https://huggingface.co/jinaai/jina-reranker-m0 to refine the initial similarity scores. This component re-evaluates the visual relevance between the target and candidate images via attention-based interactions to generate a more reliable image similarity score, $S_{\text{img}}$, and performs semantic interaction on the target and candidate texts to output a more precise text similarity score, $S_{\text{txt}}$. Finally, we fuse the scores to compute a final similarity score for each candidate sample:
(4) $S_{\text{final}}=\alpha\cdot S_{\text{img}}+(1-\alpha)\cdot S_{\text{txt}}+S_{\mathcal{T}}$
where $\alpha$ is a hyperparameter that balances the importance of the image and text modalities, and $S_{\mathcal{T}}$ represents the score derived from the structured harm matching. All candidates are sorted in descending order based on their $S_{\text{final}}$ score, from which we select the top-$K$ samples, ensuring a balanced selection by choosing half from the benign candidates and the other half from the harmful ones. The corresponding misjudgment risk patterns of these selected samples are subsequently used to guide the follow-up detection.
3.3. Pattern-Augmented Reasoning for Detection
With the retrieved memes and their associated misjudgment risk patterns, we orchestrate them into a structured prompt and guide the MLLM through a systematic reasoning chain, steering it clear of historical pitfalls to arrive at the well-justified decision.
We present pattern-augmented prompt construction, where the prompt is built upon a structured template designed to orchestrate the MLLM’s analytical workflow. The template first grounds the MLLM by clearly stating the classification task and providing an objective definition of the target harm category adopted from (Lin et al., 2025). Its core component is the Critical Reflection References section, which contains the retrieved misjudgment patterns categorized based on their ground-truth labels. Patterns associated with harmful memes are aggregated and framed as Miss Reasons to alert the MLLM to subtle harmful patterns it might otherwise overlook. Conversely, patterns from benign memes are presented as False Alarm Reasons to prevent the MLLM from misclassifying non-hateful content. To ensure this guidance is used for critical reflection rather than simple imitation, the template also enforces a structured reasoning procedure that compels the model to form an independent analysis before consulting the provided support. Through this strategy, a unique and highly relevant prompt is dynamically generated for each target sample (See Appendix C Figure 8 for the detailed prompt templates).
Finally, the complete prompt, populated with the target sample’s content and the dynamically constructed Critical Reflection References, is submitted to an MLLM. Instructed to follow the explicit reasoning procedure, the MLLM performs a context-aware analysis. This guided process ensures that the final output is not merely a binary classification but a well-reasoned judgment, critically informed by relevant historical misjudgment patterns.
4. Experiment Design
To evaluate the performance of PatMD, we introduce the following two Research Questions (RQs).
  • RQ1: What are the performances of PatMD on the harmful meme detection? We aim to evaluate the performance of PatMD against different baselines.
  • RQ2: What is the contribution of components to the harmful meme detection? We aim to analyze the effectiveness of our design, by comparing the performances of PatMD and variants of removing various components.
4.1. Datasets
We use the GOAT-bench (Lin et al., 2024), a popular comprehensive benchmark of 6,626 memes for evaluating harmful meme detection across five tasks: hateful, misogynistic, offensive, sarcastic, and harmful content. This benchmark aggregates five specialized datasets (FHM (Kiela et al., 2020), MAMI (Fersini et al., 2022), MultiOFF (Suryawanshi et al., 2020), MSD (Cai et al., 2019), Harm-C/P (Pramanick et al., 2021)). We utilize the train and test splits provided by GOAT-bench. Detailed dataset statistics are available in Table 1.
Table 1. Statistics of memes that are used in PatMD.
Type Hatefulness Misogyny Offensiveness Sarcasm Harmfulness
Train 8,500 10,000 7,000 19,816 4,250
Test 2,000 1,000 743 1,820 1,063
Table 2. Performance of our proposed PatMD and the baselines across different tasks (%).
Hatefulness Misogyny Offensiveness Sarcasm Harmfulness Overall
Approach Acc F1 Acc F1 Acc F1 Acc F1 Acc F1 Acc F1
Specialized Harmful Meme Detection Baselines
PromptHarm 64.05 51.57 69.30 67.81 61.88 57.89 58.51 55.08 61.71 53.44 63.28 57.16
Pro-Cap 52.40 51.80 61.56 58.57 48.56 47.15 54.90 54.62 57.06 56.83 54.90 53.79
CapAlign 63.70 51.60 68.60 66.81 61.51 54.76 58.08 52.67 60.77 51.04 62.53 55.38
MemeAgent 68.40 62.73 73.10 73.04 62.18 58.96 70.11 69.97 63.23 61.25 67.40 65.19
$\Delta$(GPT-4o+PatMD, SOTA) $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+12.55$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+16.41$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+13.10$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+13.14$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+2.64$} $\color[rgb]{0,0.19921875,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0}-0.02$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+12.42$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+12.44$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+5.35$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+5.44$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+9.22$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+9.48$}
MLLM-based Baselines and Our Approach
Qwen2.5VL-32B 72.85 68.25 65.20 60.89 63.17 55.33 78.24 77.95 62.26 53.88 68.34 63.26
Qwen2.5VL-32B + PatMD 74.55 72.19 78.9 78.84 63.06 60.86 81.15 81.15 65.25 63.77 72.58 71.36
$\Delta$(PatMD, Vanilla-MLLM) $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+1.70$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+3.94$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+13.70$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+17.95$} $\color[rgb]{0,0.19921875,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0}-0.11$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+5.53$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+2.91$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+3.20$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+2.99$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+9.89$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+4.24$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+8.10$}
InternVL3.5-20B-A4B 74.20 71.76 76.10 75.86 61.37 54.09 66.15 62.59 59.17 52.78 67.40 63.42
InternVL3.5-20B-A4B + PatMD 72.8 69.15 79.1 79.03 61.86 52.38 75.55 75.47 63.65 58.5 70.59 66.91
$\Delta$(PatMD, Vanilla-MLLM) $\color[rgb]{0,0.19921875,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0}-1.40$} $\color[rgb]{0,0.19921875,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0}-2.61$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+3.00$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+3.17$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+0.49$} $\color[rgb]{0,0.19921875,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0}-1.71$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+9.40$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+12.88$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+4.48$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+5.72$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+3.19$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+3.49$}
InternVL3-14B 68.25 68.24 65.90 61.82 59.08 59.06 58.68 50.34 57.86 57.75 61.95 59.44
InternVL3-14B + PatMD 76.95 75.06 77.80 77.60 63.26 57.18 80.03 79.88 66.70 64.99 72.95 70.94
$\Delta$(PatMD, Vanilla-MLLM) $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+8.70$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+6.82$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+11.90$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+15.78$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+4.18$} $\color[rgb]{0,0.19921875,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0}-1.88$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+21.35$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+29.54$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+8.84$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+7.24$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+10.99$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+11.50$}
LLaVA-1.6-13B 57.10 56.98 51.80 37.68 46.03 40.09 50.00 33.33 52.30 49.65 51.45 43.54
LLaVA-1.6-13B + PatMD 64.47 52.58 63.32 58.27 60.62 46.20 59.92 54.03 58.44 45.88 61.35 51.39
$\Delta$(PatMD, Vanilla-MLLM) $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+7.37$} $\color[rgb]{0,0.19921875,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0}-4.40$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+11.52$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+20.59$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+14.59$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+6.11$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+9.92$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+20.70$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+6.14$} $\color[rgb]{0,0.19921875,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0}-3.77$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+9.91$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+7.85$}
GPT-4o 76.60 75.24 82.71 82.43 59.16 59.13 68.85 66.27 66.32 64.09 70.73 69.43
GPT-4o + PatMD 80.95 79.14 86.2 86.18 64.82 58.94 82.53 82.41 68.58 66.69 76.62 74.67
$\Delta$(PatMD, Vanilla-MLLM) $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+4.35$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+3.90$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+3.49$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+3.75$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+5.66$} $\color[rgb]{0,0.19921875,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0}-0.19$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+13.68$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+16.14$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+2.26$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+2.60$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+5.89$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+5.24$}
GPT-4o-mini 68.85 66.41 79.30 78.92 57.95 57.82 57.97 49.29 66.50 66.10 66.11 63.71
GPT-4o-mini + PatMD 71.41 66.97 78.70 78.30 64.62 62.47 80.27 80.10 68.02 67.64 72.60 71.10
$\Delta$(PatMD, Vanilla-MLLM) $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+2.56$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+0.56$} $\color[rgb]{0,0.19921875,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0}-0.60$} $\color[rgb]{0,0.19921875,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0}-0.62$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+6.67$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+4.65$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+22.30$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+30.81$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+1.52$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+1.54$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+6.49$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+7.39$}
4.2. Baselines
We compare our approach against two categories of baselines: specialized approaches for harmful meme detection and general-purpose MLLMs.
Specialized Approaches for Harmful Meme Detection. This category includes methods specifically designed or fine-tuned for this task. We consider two sub-types. The first are supervised methods that require task-specific training, including PromptHarm (Cao et al., 2022), Pro-Cap (Cao et al., 2023), and CapAlign (Ji et al., 2024). The second, which we refer to as MemeAgent (Lin et al., 2025), is a SOTA multi-agent framework for meme analysis that operates without requiring end-to-end training. For a fair comparison, we utilize the implementations of all these methods as provided in the MemeAgent study.
General-Purpose MLLMs. To demonstrate the effectiveness and generalizability of our approach, we establish baselines using a range of SOTA, general-purpose MLLMs. This includes Qwen2.5-VL-32B-Instruct (Bai et al., 2025), InternVL3.5-20B-A4B (Wang et al., 2025), InternVL3-14B (Zhu et al., 2025), LLaVA-1.6-13B [^3] The baseline performance for LLaVA-1.6-13B is sourced from the MemeAgent (Lin et al., 2025) study for consistency.  (Liu et al., 2024), GPT-4o (Hurst et al., 2024), and GPT-4o-mini (Hurst et al., 2024). We evaluate their native performance in a zero-shot CoT setting.
4.3. Metrics
To ensure a comprehensive and impartial evaluation, we measure model performance using two core metrics: Accuracy and Macro-F1 Score. This choice aligns with the established evaluation protocol of the GOAT-Bench, thereby enabling a direct and consistent comparison with prior work.
4.4. Experimental Setup
All experiments were conducted on a server equipped with a single NVIDIA H100 GPU with 80GB of VRAM. For the open-source models, we deployed local inference endpoints using the vLLM framework (Kwon et al., 2023), loading pre-trained weights sourced from the HuggingFace Hub. For proprietary models, we accessed them via their official APIs, specifically using GPT-4o (gpt-4o-2024-11-20 version) and GPT-4o-mini (gpt-4o-mini-2024-07-18 version). Specifically, for the Misjudgment Risk Pattern Elicitation stage, we utilized GPT-4o as the MLLM.
To improve reproducibility, the temperature parameter was set to 0 during inference, which corresponds to deterministic decoding without sampling. In addition, the prompting strategy for the baseline strictly followed the configuration described in the previous work (Lin et al., 2025) to guarantee fairness and consistency in comparison.
The hyperparameters for our method are set as follows: top-NN=10, top-KK=2, $\alpha=0.7$, $\beta=0.7$, and $\gamma=0.1$.
5. Results and Analysis
Table 3. The results of ablation studies of the PatMD (%).
Approach Hatefulness Misogyny Offensiveness Sarcasm Harmfulness Overall
Acc F1 Acc F1 Acc F1 Acc F1 Acc F1 Acc F1
PatMD 76.95 75.06 77.80 77.60 63.26 57.18 80.03 79.88 66.70 64.99 72.95 70.94
w/o tree_matching 74.26 71.36 77.30 77.19 62.25 55.37 78.26 78.18 64.11 62.22 71.24(-1.71) 68.86(-2.08)
w/o img_txt_rerank 76.19 74.29 76.98 76.73 62.05 56.32 78.95 78.78 66.13 64.49 72.06(-0.89) 70.12(-0.82)
w/o txt_retrieval 75.70 73.74 77.34 77.29 62.48 56.39 78.04 77.89 64.35 62.26 71.58(-1.37) 69.51(-1.43)
w/o img_retrieval 75.64 73.67 77.10 76.83 62.99 55.77 79.35 79.20 65.10 62.99 72.04(-0.91) 69.69(-1.25)
random retrieval 73.07 69.83 74.54 74.54 61.89 54.15 75.06 74.80 63.55 61.00 69.62(-3.33) 66.86(-4.08)
5.1. Performance of PatMD (RQ1)
To evaluate the effectiveness of our proposed approach, we compare the performance of PatMD with the baselines and general-purpose MLLMs (i.e., Vanilla-MLLM in the table), as shown in Table 2.
Overall Performance. The results demonstrate that PatMD consistently improves the performance of specialized baselines and general-purpose MLLMs, with an average of 8.30% improvement in F1-scores and an average of 7.71% improvement in accuracy. First, when compared against specialized harmful meme detection baselines, our method shows strong performance. For instance, when applied to GPT-4o and the open-source InternVL3-14B, our framework achieves overall F1-scores of 74.67% and 70.94% respectively, both outperforming the strongest baseline, MemeAgent (65.19%). Second, the framework exhibits strong generalization across different MLLMs. It improves the overall F1-score of GPT-4o by 5.24% and that of GPT-4o-mini by 7.39%. The impact on open-source models is also notable. InternVL3-14B experiences an 11.50% F1-score increase, surpassing the vanilla GPT-4o (70.94% vs. 69.43%). Similarly, LLaVA-1.6-13B, a model of comparable scale, receives an F1-score boost of 7.85%. Furthermore, the framework proves effective even for specialized architectures such as InternVL3.5-20B-A4B, a Mixture-of-Experts (MoE) model activating only 4B parameters during inference, achieving an overall F1 improvement of 3.49%. These results validate that PatMD is an effective and model-agnostic approach capable of enhancing the detection performance of MLLMs across different scales and architectures.
Task-Specific Analysis. A detailed analysis of task-specific results further reveals the breadth and depth of the improvements. The framework’s impact is particularly evident in Sarcasm and Misogyny detection. Notably, for Sarcasm, PatMD yields substantial improvements across models. For instance, InternVL3-14B’s F1-score improves by 29.54%, and GPT-4o-mini’s F1-score increases by 30.81%. This highlights PatMD’s ability to help models interpret the incongruity between text and image—a common characteristic of sarcastic memes. In Misogyny detection, similar improvements are observed. Qwen2.5VL-32B’s F1-score, for example, increases by 17.95%. This indicates that by retrieving historical samples with similar stereotypical patterns, our framework effectively equips models to identify targeted, gender-based harm. Beyond these tasks, PatMD also demonstrates effectiveness across other domains. In Offensiveness detection, most models show positive gains, such as the 6.67% F1-score increase for GPT-4o-mini. For Harmfulness, a general positive trend is also observed, with Qwen2.5VL-32B gaining 9.89% in F1-score. These widespread gains underscore the framework’s ability to provide task-relevant reasoning support.
5.2. Ablation Study (RQ2)
To evaluate the effectiveness of the key components in our PatMD, we conduct a series of ablation studies. We evaluate the impact on overall performance by removing or replacing key components of our approach, including: the structural matching of hierarchical harm tree in Section 3.2.2 (w/o tree_matching), the image and text reranking in Section 3.2.2 (w/o img_txt_rerank), the text/image retrieval in Section 3.2.1 (w/o txt/img_retrieval), and replacing the bipartite candidate retrieval of Section 3.2.1 with random retrieval (random retrieval). The experimental results are presented in Table 3.
As shown in Table 3, the complete PatMD approach achieves the best performance in both accuracy (Acc) and F1 score, demonstrating the integrity and synergistic effect of our design. Among all variants, removing the structural matching of hierarchical harm tree (w/o tree_matching) leads to one of the most significant performance drops, with the F1 score decreasing by 2.08% (from 70.94% to 68.86%). This strongly demonstrates that our proposed hierarchical harm tree is crucial for the model to understand and judge the deep-seated risks in complex memes. Removing the image and text reranker (w/o img_txt_rerank), which is used to refine content similarity, also results in a performance decline. This indicates that a more precise assessment of content relevance is essential besides performing the structure matching. Furthermore, performance drops when the text retrieval (w/o txt_retrieval) and image retrieval (w/o img_retrieval) components are removed individually, validating the necessity of leveraging both modalities for retrieval, as relying on a single modality can miss key similar samples. Finally, compared to random retrieval (random retrieval), our method shows a substantial performance advantage, with a F1 score more than 4% higher. This highlights that accurately retrieving highly relevant historical memes from the knowledge base is vital to ensuring the overall effectiveness of the framework.
In addition, we further evaluate the impact of the number of retrieved samples ($K$) on PatMD’s performances, and the results are shown in the Appendix A.
6. Discussion
Figure 4. The case study of PatMD on four examples.
In this section, we provide several cases to intuitively illustrate the advantages of PatMD with the misjudgment risk patterns, as well as the human evaluation of the pattern’s quality.
6.1. Case Study
To more intuitively understand our framework’s reasoning process, we first analyze several correctly predicted cases. These cases demonstrate the misjudgment risk patterns retrieved by our approach and show the key content of the reasoning process. We also conduct a failure case analysis to explore the current limitations and future directions of our approach, as shown in Figure 4.
Detecting Implicit Harm Masked by Neutral Language and Humor. A significant challenge in meme detection is identifying harmful intent hidden behind seemingly innocuous text or humor. Our approach excels in this area by retrieving relevant risk patterns that guide MLLM to look beyond the surface. In Sample (a), the meme uses a common phrase, “spot the difference.” Our approach, guided by the retrieved misjudgment risk pattern warning against “neutral text paired with images that implicitly dehumanize,” correctly identifies the racist comparison. The output reasoning clearly states that “the juxtaposition of monkeys and children carries an implicit risk of dehumanization,” leading to the correct “Hateful” prediction. Similarly, in Sample (b), the meme employs irony to convey a misogynistic message. Our approach retrieves a pattern highlighting that “humor or irony can obscure misogyny.” This prompts a deeper analysis, as reflected in the output: MLLM correctly discerns that the meme “trivializes women’s rights by equating equality with physical confrontation,” thus identifying its underlying misogynistic nature.
Accurate Sarcasm Detection through Contextual Incongruity. Our approach’s ability to analyze context extends to non-harmful categories like sarcasm, where understanding the intent is key. The task for Sample (c) is to identify if the content is sarcastic. The meme uses an absurd analogy of a panda to deliver an anti-racist message. Our system retrieves risk patterns related to “irony depends on sociopolitical context” and situations where “text…contradicts absurd imagery.” Guided by these insights, MLLM correctly focuses on the incongruity between the serious issue of racism and the simplistic panda analogy. This understanding allows it to correctly classify the meme as “Sarcastic,” demonstrating the approach’s ability to interpret nuanced forms of expression.
Failure Analysis: The Subjective Boundary Between Satire and Offense. In Sample (d), a political meme satirizing Donald Trump is labeled “Offensive,” but our model predicted “Benign.” The analysis reveals a key challenge: while the approach successfully retrieved the relevant risk pattern— “When offensive content is framed through satire…it may be overlooked”—it struggled with the final judgment. It over-prioritized the satirical context and the targeting of a public figure, leading to an incorrect classification. This highlights the difficulty in navigating the subjective boundary between sharp satire and offensive content, marking a clear area for future improvement.
In summary, these case studies demonstrate that our misjudgment driven approach enhances a model’s ability to decipher the implicit and contextual cues in memes. This leads to a more nuanced understanding and accurate classification across different categories, from hate speech to sarcasm.
Figure 5. Distribution of human evaluation scores for misjudgment risk patterns generated by GPT-4o.
6.2. Human Evaluation
To validate the quality of the misjudgment risk patterns generated by GPT-4o, we conduct human evaluation by referencing the evaluation framework proposed by (Mazhar et al., 2025). We randomly sample 20 instances from each of the five tasks within our benchmark, creating an evaluation set of 100 instances in total. Each meme is associated with a generated misjudgment risk pattern.
We instruct three human evaluators, each with over five years of experience in digital content moderation, to rate these patterns on a 5-point Likert scale (from 1: poor to 5: excellent) based on four key criteria: relevance, correctness, actionability, and uniqueness. The evaluation results, as illustrated in Figure 5, clearly indicate that the majority of ratings are concentrated between 4 and 5. This suggests that the misjudgment risk patterns generated by GPT-4o are generally well-perceived and align closely with human judgment on potential risks.
7. Conclusion
In this paper, we propose PatMD, an approach that enhances harmful meme detection by learning from potential misjudgment risks. It first constructs a knowledge base, where each meme is deconstructed into a misjudgment risk pattern that captures why and how the MLLM may misinterpret its content. For a given target meme, PatMD retrieves the semantically similar memes, and the relevant patterns are used to dynamically guide the MLLM’s reasoning, proactively guiding the MLLMs to avoid known misjudgment pitfalls. Experiments show that PatMD significantly outperforms the baselines, achieving an average of 7.26% F1-score improvement.
Appendix A Analysis of Retrieved Samples ($K$)
Figure 6. Impact of the retrieved case’s number ($K$).
In addition to the ablation study in the Section 5.2, we also investigated the impact of the number of retrieved samples (top-$K$) on model performance, with the results shown in Figure 6.
As illustrated in the figure, as the value of $K$ increases, the model’s performance on most tasks shows a trend of initially rising and then stabilizing or slightly declining. Specifically, the model achieves optimal or near-optimal performance on most tasks when $K$ is between 2 and 4. This suggests that providing a small number of high-quality misjudgment risk patterns from similar memes can effectively guide the model’s reasoning. However, when $K$ becomes too large (e.g., greater than 4), introducing too much information may introduce noise that interferes with the model’s judgment. Considering both performance and efficiency, we set the optimal value of $K$ to 2 in our experiments.
Appendix B Pseudocode for PatMD
In this section, we have provided the pseudocode of the pattern retrieval, so as to improve the reproducibility.
Input: Test sample $m_{test}$, Knowledge base $K$.
Output: Final prediction $y_{pred}$ and rationale analysis.
1 end Function RetrieveAndDetect($m_{test},K$):
2  $v_{test}^{\text{img}},v_{test}^{\text{txt}}\leftarrow E(\text{img}_{test},\text{txt}_{test})$;
3  $\text{Candidates}_{b}\leftarrow\textsc{Retrieve}(K,\text{'benign'},v_{test}^{\text{img}},v_{test}^{\text{txt}},N)$;
4  $\text{Candidates}_{h}\leftarrow\textsc{Retrieve}(K,\text{'harmful'},v_{test}^{\text{img}},v_{test}^{\text{txt}},N)$;
5  $\text{CandidatePool}\leftarrow\text{Candidates}_{b}\cup\text{Candidates}_{h}$;
6
7 $\text{ScoredList}\leftarrow[]$;
8  for each candidate $c$ in CandidatePool do
9     $S_{\text{img}}\leftarrow R_{img}(\text{img}_{test},c.\text{img})$;
10     $S_{\text{txt}}\leftarrow R_{txt}(\text{txt}_{test},c.\text{txt})$;
11
12    $S_{\mathcal{T}}\leftarrow\textsc{MatchKnowledgeTree}(m_{test}.\text{tree},c.\text{tree})$;
13
14    $S_{\text{final}}\leftarrow\alpha\cdot S_{\text{img}}+(1-\alpha)\cdot S_{\text{txt}}+S_{\mathcal{T}}$;
15     Append $(c,S_{\text{final}})$ to ScoredList;
16  end for
17 Sort ScoredList in descending order by $S_{\text{final}}$;
18  $\text{TopK\_Cases}\leftarrow$ Select top $K$ items from ScoredList;
19
20 $P_{Hateful}\leftarrow\textsc{ExtractPattern}(\text{TopK\_Cases},\text{'hateful'})$;
21  $P_{Benign}\leftarrow\textsc{ExtractPattern}(\text{TopK\_Cases},\text{'benign'})$;
22  $\text{prompt}\leftarrow\textsc{ConstructPrompt}(m_{test},P_{Hateful},P_{Benign})$;
23  $(\text{analysis},y_{pred})\leftarrow\text{MLLM}_{\text{judge}}.\textsc{Query}(\text{prompt})$;
24
25
return analysis, $y_{pred}$
Algorithm 1 Risk-aware Pattern Retrieval.
Appendix C Prompt for Harmful Meme Detection
To improve the reproducibility of PatMD, we provide the contents of three prompts that are used in the harmful meme detection:
  • Figure 7’s Prompt: Hierarchical harm tree construction (Step-1 of PatMD), where the prompt is used in Section 3.1.1 hierarchical harm deconstruction.
  • Figure 8’s Prompt: Harmful meme detection (Step-3 of PatMD), where the prompt is used in Section 3.3 pattern-augmented prompt construction.
  • Figure 9’s Prompt: Misjudgment risk pattern generation (Step-1 of PatMD), where the prompt is used in Section 3.1.2 causal misjudgment attribution.
Figure 7. Prompt for hierarchical harm tree construction.
Figure 8. Prompt for harmful memes detection.
Figure 9. Prompt for misjudgment risk pattern generation.